OpenStudy (anonymous):

Pretty please help me !! What is the slope of this graph when it is changing most rapidly ? $A\left(x\right)=\frac{43}{1+64.4e^{\left(-0.79x\right)}}$

OpenStudy (anonymous):

\[A\left(x\right)=\frac{43}{1+64.4e^{\left(-0.79x\right)}}\]

OpenStudy (anonymous):

A(x)=\[\frac{43}{1+64.4e^{\left(-0.79x\right)}}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i am not sure what "slope of the graph when it is changing the most" means

OpenStudy (anonymous):

most rapidly

OpenStudy (anonymous):

maybe take the second derivative to find when the first derivative is largest

OpenStudy (anonymous):

here is a picture, lets see if we can find the second derivative http://www.wolframalpha.com/input/?i= \frac{43}{1%2B64.4e^{\left%28-0.79x\right%29}}+domain+-2..10

OpenStudy (anonymous):

wow this is going to suck to do by hand here is the derivative, and below is the second derivative http://www.wolframalpha.com/input/?i=%282187.67+e^%280.79+x%29%29%2F%2864.4%2Be^%280.79+x%29%29^2&lk=1&a=ClashPrefs_*Math-

OpenStudy (anonymous):

looks like the derivative has a local max at \(5.2723\) so that is the \(x\) value where the function is changing most rapidly

OpenStudy (anonymous):

Okay thank you !!

OpenStudy (anonymous):

yw, hope it is right